OpenStudy (anonymous):

An airline has 81% of its flights depart on schedule. It has 69% of its flights depart and arrive on schedule. Find the probability that a flight that departs on schedule also arrives on schedule. 1.25 0.85 0.45 0.43

OpenStudy (anonymous):

probablity=desired divided by total I don't know if this could help?

OpenStudy (anonymous):

uhh Idk, I need the steps :P @Mertsj help please?

OpenStudy (anonymous):

isnt that the 69%?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

well it says 69% departs and arrives on time and then the question asks the probability of flights that depart and land on schedule

OpenStudy (anonymous):

so wouldnt you do 69 divided by 81?

OpenStudy (anonymous):

Ummmmmm... I wish I knew, I guess that could be right ? :P Do you know for sure that's what to do?

OpenStudy (anonymous):

Sorry, i'm just terrible at statistics :P

OpenStudy (mertsj):

A = departs on schedule P(A)=0.81 B = arrives on schedule P(A and B) = 0.69 P(B/A) = P( A and B) / P(A) = 0.69/0.81 = 0.85

OpenStudy (anonymous):

yeah thats what you do...because it is giving you the starting percent of 81 being on time...out of that 69 has both so you would divide 69/81

OpenStudy (anonymous):

Ohhh, thanks soo much guys, I would give medal to both but I'll give to Hope since she needs more :)

OpenStudy (anonymous):

well thanks :D

OpenStudy (mertsj):

Me too.

OpenStudy (anonymous):

awh you guys are awesome :D